OpenStudy (chris215):

what is the equilibrium temperature and why is the change in temperature for A and B the same? Initial temp. A: 100 Initial temp. B: 0 Final temp. A: 50 Final temp. B: 50 Change in temp. A: -50 Change in temp. B: 50

OpenStudy (ybarrap):

What is the final temperature of A? What is the final temperature of B? These are the equilibrium temperatures of each You have the change above, one of them gets colder, the second gets warmer. The magnitudes of the change is the same, but the direction of these changes is different.

OpenStudy (chris215):

the specfic heat capacity for A and B is 387

OpenStudy (ybarrap):

$$ C \equiv \frac{ Q}{\Delta T} $$ Since they have the same heat capacity $$ 387=\frac{ Q_A}{\Delta T_A}=\frac{ Q_B}{\Delta T_B}\\ $$ Which means \(Q_B=-Q_B\) Do you agree?

OpenStudy (chris215):

yep thanks!